The standard dose of medicine for a dog is 10 grams for every15 pounds of body weight. Whatis the standard dose for a dog that weighs 80 pounds?
Answer by checkley77(12569)

15/80=10/x
15x=10*80
15x=800
x=800/15
x=53 1/3 grams is needed.